In a recent awards​ ceremony, the age of the winner for best actor was 33 and the age of the winner for best actress was 59. For all best​ actors, the mean age is 42.8 years and the standard deviation is 8.3 years. For all best​ actresses, the mean age is 35.9 years and the standard deviation is 12.1 years.​ (All ages are determined at the time of the awards​ ceremony.) Relative to their​ genders, who had the more extreme age when winning the​ award, the actor or the​ actress? Explain.


Since the z score for the actor is z=

nothing


and the z score for the actress is z=

nothing


​, the ▼

 

actor

actress

had the more extreme age.

​(Round to two decimal​ places.)


1
Expert's answer
2021-09-07T13:30:38-0400

"Z_{actor}=\frac{33-42.8}{8.3}= -1.18 \\

Z_{actress}= \frac{59-35.9}{12.1}= 1.91"

Since the z score for the actor is z= -1.18 and the z score for the actress is z= 1.91, the actress had the more extreme age.